OpenStudy (anonymous):

Katie invested $1200 at 8.5% interest. She earned $612 interest at the end of y years. What is the value of y? A. 18 B. 12 C. 6 D. 3

OpenStudy (cggurumanjunath):

(si)=ptr/100

OpenStudy (cggurumanjunath):

si = simple interest

OpenStudy (cggurumanjunath):

612=si

OpenStudy (cggurumanjunath):

p=1200

OpenStudy (cggurumanjunath):

r=8.5

OpenStudy (cggurumanjunath):

(si)=ptr/100 612=1200*t*8.5/100

OpenStudy (cggurumanjunath):

now can u solve for t @Kmcerullo

OpenStudy (cggurumanjunath):

t=((612*100)/(1200*8.5))

OpenStudy (anonymous):

Im confused

OpenStudy (cggurumanjunath):

how much is 612*100

OpenStudy (cggurumanjunath):

how much is 1200*8.5

OpenStudy (anonymous):

Okay i got it thank you :)

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

Lily invested $750 in a savings account and earned $225 in interest at the end of 5 years. What was the interest rate? A. 5% B. 5.5% C. 6% D. 6.5% @CGGURUMANJUNATH Can you help me ?

OpenStudy (cggurumanjunath):

(si)=ptr/100

OpenStudy (cggurumanjunath):

now u solve this prob i will help u !

OpenStudy (cggurumanjunath):

si=?

OpenStudy (anonymous):

i just need the steps for this one and ill pro get it

OpenStudy (cggurumanjunath):

that is what i am helping u

OpenStudy (cggurumanjunath):

si=225 right ?

OpenStudy (cggurumanjunath):

now u tell me p=?

OpenStudy (cggurumanjunath):

t=?

OpenStudy (anonymous):

p=750 ?

OpenStudy (cggurumanjunath):

yes

OpenStudy (cggurumanjunath):

t=?

OpenStudy (anonymous):

and t=5 ?

OpenStudy (cggurumanjunath):

yes

OpenStudy (cggurumanjunath):

(si)=ptr/100

OpenStudy (cggurumanjunath):

r has to b found out !

OpenStudy (anonymous):

how do i find what r is ?

OpenStudy (cggurumanjunath):

by substituting in the formula (si)=ptr/100

OpenStudy (anonymous):

thats confusing lol

OpenStudy (cggurumanjunath):

225=(750*5*r)/100

OpenStudy (cggurumanjunath):

(225*100)=(750*5*r)

OpenStudy (anonymous):

thank you :)

OpenStudy (cggurumanjunath):

(225*100)/(750*5)=r

OpenStudy (cggurumanjunath):

r=? @Kmcerullo

OpenStudy (cggurumanjunath):

tell me the final answer !

OpenStudy (anonymous):

6

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

u got it .

OpenStudy (anonymous):

Thank you

OpenStudy (cggurumanjunath):

WC.